OpenStudy (anonymous):

can someone check if this equation is true please? 2√(7.5+5)+7=4

OpenStudy (anonymous):

true

OpenStudy (anonymous):

can you show me your work?

OpenStudy (anonymous):

its not true btw

OpenStudy (anonymous):

i just don't know how to do it

OpenStudy (anonymous):

but ik its not true

jigglypuff314 (jigglypuff314):

first try the sqrt(7.5+5) I feel like that would turn into a really weird number xD

OpenStudy (anonymous):

12.5

OpenStudy (anonymous):

i did up to that but i get confused when to get rid of the square root

OpenStudy (phi):

are you allowed to use a calculator?

OpenStudy (anonymous):

yes

OpenStudy (phi):

then do 2* sqrt(12.5)+7

OpenStudy (anonymous):

14.0710678119

OpenStudy (phi):

but you should know that 3*3 is 9 and 4*4 is 16 so the square root is between 3 and 4 and if we just use 3 2*3+7= 13 and if we used 4 it would be even bigger. we will never get 4

OpenStudy (phi):

normally an equation has a variable (a letter) are you sure you have the correct question ?

OpenStudy (anonymous):

the question was to make an equation with an extraneous solution, i did 2√(x+5)+7=4 and it worked out to 7.5, now I'm checking if its extraneous

OpenStudy (phi):

2√(7.5+5)+7=4 14.07... \( \ne\) 4

OpenStudy (anonymous):

thats it?

OpenStudy (phi):

so you solved for x ? 2√(x+5)+7=4 \[ \sqrt{x+5}= -\frac{3}{2} \\ x+5 = \frac{9}{4} \\ x= -\frac{11}{4} \]

OpenStudy (phi):

that solution is extraneous (it does not work in the original equation)

OpenStudy (anonymous):

Thank you sir

OpenStudy (anonymous):

you are a very good teacher